THE ROTATIONAL SPECTRUM OF CHLORINE NITRATE (ClONO_{2}) IN THE THREE LOWEST nν_{9} POLYADS

\underline{Z. KISIEL}, E. BIAŁKOWSKA-JAWORSKA, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland; R. A. H. BUTLER, Department of Physics, Pittsburg State University, Pittsburg, KS 66726; D. T. PETKIE, Department of Physics, Wright State University, Dayton, OH 45435; P. HELMINGER, Department of Physics, University of South Alabama, Mobile, AL 36688; F. C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210.

The rotational spectrum of the stratospherically important ClONO_{2} molecule has recently been subjected to an extended analysis,^{a} which covered four states ranging from the ground state to the ν_{5}/ν_{6}ν_{9} dyad at ≈560 cm^{−1}. The analysis was based on a newly recorded broadband FASSST spectrum with an almost complete coverage of the 118-378 GHz region.

The remaining vibrational states below 560 cm^{−1} are the three polyads associated with successively higher excitation of the ν_{9} mode. These are the (2ν_{9}/ν_{7}) dyad, the (3ν_{9}/ν_{7}ν_{9}) dyad, and the (4ν_{9}/ν_{7}2ν_{9}/2ν_{7}) triad. Rotational transitions in the triad were assigned for the first time and data sets for the two dyads were considerably improved over those obtained previously.^{b} Transition frequencies in all three polyads, for both ^{35}ClONO_{2} and ^{37}ClONO_{2}, were fitted to within experimental accuracy by using a new coupling scheme which accounted in a unified manner for the pertinent Fermi and Coriolis interactions. Various tests used to ensure that the ’true’ solution has been reached are discussed.

aR.A.H.Butler et al., J. Mol. Spectrosc., 243, 1-9 (2007); 244, 113-116 (2007).

bR.A.H.Butler et al., J. Mol. Spectrosc., 213, 8-14 (2002); 220, 150-152 (2003).